OpenStudy (anonymous):

The ratio of the numerator to the denominator of a certain fraction is one to four. If three is added to the numerator and subtracted from the denominator, the new fraction reduces to one-third. What is the original fraction? 12/45 15/45 12/48 i got b

OpenStudy (anonymous):

let fraction is \[\frac{ x }{ y }\] \[then ~\frac{ x }{y } =\frac{ 1 }{4 },~or~y=4x\] again \[\frac{ x+3 }{y-3 }=\frac{ 1 }{3 },~or~3x+9=y-3\] substitute the value of y, then find x and finally y.

OpenStudy (anonymous):

it is not b